# Machine-learning for photoplethysmography analysis: Benchmarking feature, image, and signal-based approaches

### v1

@@ 3.3 Prediction models
In our study, we explore the use of three different input representations —time series, feature-based, and image-based approaches—each providing unique advantages in capturing the underlying physiological information. We consider several architectures, each designed to process either raw signals, extracted features, or visual representations to accurately estimate BP values and classify heart rhythms.
Baseline Models. On VitalDB CalibFree, the baseline model (used for BP estimation) predicts BP by outputting the median SBP/DBP value of the blood pressure data inferred from the training set for any given input, providing a straightforward reference for evaluating the performance of more advanced predictive models. On VitalDB Calib, we use the subject-specific median calculated as prediction on the test set.
Raw Time Series Models. For both BP estimation and AF detection, deep learning architectures such as CNNs, RNNs, and TCNs have been used to process raw ECG or PPG sequences to capture complex temporal patterns, directly predicting continuous values for regression or output probabilities for classification. The specific models used in this study were:
@@
  • XResNet1d50/101: Deep residual networks modified for 1D data, incorporating group normalization and selective kernel sizes to learn hierarchical features from physiological signals. [34].
  • XResNet1d50+GNNLL: This model explores the XResNet50d with a Gaussian Negative Log Likelihood Loss (GNLL) instead of the conventional MAE loss as proposed in [35].
@@
We can roughly categorize the considered models into complex/deep models (Inception1D, XResNet1d50, XResNet1d101) and simple/shallow models (LeNet1d, AlexNet1d, MiniRocket, TCNs).
Feature-Based Models. Clinically interpretable features are extracted from PPG signals—such as pulse morphology metrics for BP estimation and irregularity measures for AF detection—and fed into machine learning algorithms to perform regression or classification tasks with clearer interpretability. The specific models and techniques used in this study were:
  • Clinically interpretable features (CIF): CIF for BP includes features such as systolic peaks, diastolic peaks, pulse duration, and pulse morphology metrics [39]. These features reflect vascular health, haemodynamic dynamics, and arterial compliance, making them well-suited for modeling BP. CIF for AF comprises features related to rhythm irregularity, such as randomness [40], variability [41, 42], and complexity in inter-beat-intervals [43]. These features highlight the irregular heart rhythms characteristic of atrial fibrillation. The full list of features is described in the supplementary material. These features are combined with the following prediction models:
@@
  • Wavelet Transformation: A Wavelet-based method such as Wavelet Packet Decomposition using the Discrete Wavelet Transform [45] is applied to the raw PPG signal. This transformation decomposes the signal into time-frequency components, extracting features that capture both transient and long-term patterns in the data. It should be noted that Wavelet + Multi-Layer Perceptron (MLP) forms a pipeline where the wavelet transform serves as a sophisticated feature extractor, and the MLP functions as the predictive model utilizing those features.
Image-Based Models: One-dimensional signals can be converted into two-dimensional images which capture the essence of the signal in a compact domain. Traditional image recognition CNNs can then be used with these image inputs. There are different ways of converting signals to images, but this study only used the following approach:
  • Continuous Wavelet Transform (CWT) Scalograms: PPG signals are transformed into time-frequency images derived from PPG signals that capture localized signal variations (26). The CWT-based images are used as inputs for ResNet18 models.
For more details on these models and their implementations, please see the Supplementary Material.
@@ 4.1 Blood Pressure Estimation
Tables LABEL:table_3LABEL:table_4 display the results of blood pressure (SBP/DBP) prediction using different deep learning models based on VitalDB Calib and VitalDB CalibFree datasets, respectively. The baseline models achieved MAEs of 14.87 and 9.43 mmHg for SBP and DBP respectively on CalibFree, and 10.72 and 5.78 mmHg respectively on Calib. On Calib, the per-subject baseline performed substantially better than the global baseline (10.72 and 5.78 mmHg vs. 14.91 and 9.52 mmHg respectively).
@@
Model performance varied between models and tasks. On CalibFree, almost all models provided an improvement over baseline for both SBP and DBP (the only exception being the Inception1d SBP model). However, the level of improvement was moderate at best, with the lowest MASEs of 0.83 for both SBP and DBP achieved by XResNet1d50, indicating 17% reductions in MAE in comparison to baseline. This resulted in at best 26% of SBP estimates and 40% of DBP estimates falling into the top grade (i.e. errors of <5 mmHg). On Calib, there was greater variation in model performances, with XResNet1d50+GNLL achieving the lowest MASEs of 0.73 and 0.87 for SBP and DBP respectively, corresponding to 48% of SBP estimates and 64% of DBP estimates falling into the top grade. In contrast, several models performed worse than the per-subject baseline, with the worst-performing model, CIF+MLP, achieving MASEs of 1.27 and 1.54 for SBP and DBP respectively. Indeed, for DBP estimation only the XResNet1d50+GNLL model achieved an improvement over the subject-specific baseline, whereas all others performed worse than this baseline. Absolute performance in terms of MAE was better on Calib than CalibFree, as shown by lower MAEs on Calib than CalibFree for all models except the TCN+MLP SBP and DBP models, and the CIF+MLP DBP model. MLP models always performed worse than other models of the same type.
@@
Table 3: The performance analysis for the regression task on the VitalDB CalibFree dataset for three input representations, T for raw time series, F for feature-based, and I for image-based models, next to B for the baseline model. The best-performing model is marked in bold-face and underlined, while the second and third best-performing models are highlighted in bold-face for each subset. All MAE values are given in units of mmHg.
Model SBP MAE IEEE Grades for SBP DBP MAE IEEE Grades for DBP
(MASE) A B C D (MASE) A B C D
B Baseline 14.87 (1.00) 0.21 0.04 0.04 0.71 9.43 (1.00) 0.33 0.06 0.06 0.55
T XResNet1d101 12.70 (0.85) 0.25 0.05 0.05 0.65 8.05 (0.85) 0.39 0.07 0.06 0.48
XResNet1d50 12.40 (0.83) 0.24 0.05 0.05 0.66 7.84 (0.83) 0.40 0.07 0.06 0.47
Inception1d 14.97 (1.01) 0.21 0.04 0.04 0.71 8.98 (0.95) 0.28 0.05 0.05 0.62
LeNet1d 12.37 (0.83) 0.25 0.05 0.05 0.65 7.89 (0.84) 0.39 0.07 0.06 0.47
XResNet1d50+GNLL 12.48 (0.84) 0.26 0.05 0.04 0.65 8.16 (0.87) 0.38 0.07 0.06 0.49
Alexnet1d 12.34 (0.83) 0.25 0.05 0.05 0.65 7.88 (0.84) 0.39 0.07 0.07 0.47
Minirocket 12.35 (0.83) 0.26 0.05 0.05 0.64 7.91 (0.84) 0.39 0.07 0.07 0.47
TCN +MLP 12.72 (0.85) 0.25 0.05 0.05 0.65 8.24 (0.87) 0.38 0.06 0.06 0.49
F WAVELET +MLP 14.21 (0.95) 0.22 0.04 0.04 0.70 8.89 (0.94) 0.35 0.06 0.06 0.53
CIF +GPR 12.90 (0.87) 0.25 0.05 0.05 0.65 8.15 (0.86) 0.38 0.07 0.07 0.48
CIF +MLP 14.02 (0.94) 0.24 0.04 0.04 0.68 8.69 (0.92) 0.36 0.06 0.06 0.52
I CWT 13.40 (0.90) 0.24 0.05 0.05 0.66 8.39 (0.88) 0.37 0.07 0.07 0.49
@@
Raw time series (T) and image-based (I) models performed better than feature-based (F) models on Calib (with the one exception of TCN+MLP), as shown by MASEs of 0.73-1.07 (SBP) and 0.87-1.46 (DBP) for raw time series and image-based models, compared to 1.13-1.27 and 1.35-1.54 for feature-based models. There was a less clear difference in performance on CalibFree. Whilst best performance was achieved with raw time series models, the image-based model performed better than three raw time series models on Calib, and better than one raw time series model on CalibFree. For Calib, more complex models seem to exhibit an advantage (comparing for example XResNet1d50 and XResNet1d101) most likely due to the ability of memorizing subject-specific signal patterns. The only exception from this pattern seems to be the AlexNet1d model, which shows a performance that is almost on par with more complex models and considerably better than comparable lightweight models (such as LeNet1d or TCN+MLP).
@@
Absolute performance in terms of MAE was always better for DBP estimation than SBP estimation. However, when considering errors relative to baseline, MASEs were broadly similar between SBP and DBP on CalibFree, and always higher for DBP than SBP on Calib. Since the IEEE grading system uses absolute errors, performance in terms of IEEE Grades was generally better for DBP than SBP.
4.2 Atrial Fibrillation Detection
Table 4: The performance analysis for the classification task on the DeepBeat dataset for three input representations: T for raw time series, F for feature-based, and I for image-based models. The best-performing model is marked in bold-face and underlined, while the second and third best-performing models are highlighted in bold-face for each subset.

### v2

@@ 3.3 Prediction models
In our study, we explore the use of three different input representations – time series, feature-based, and image-based approaches – each providing unique advantages in capturing the underlying physiological information. We consider several architectures, each designed to process either raw signals, extracted features, or visual representations to accurately estimate BP values and classify heart rhythms. All models were trained from scratch. The learning rate, as the most important hyperparameter, was systematically optimized based on the validation set performance, while model‑dependent hyperparameters were deliberately not extensively tuned to preserve a balanced experimental setting. Due to the large number of considered model architectures, a more systematic exploration of hyperparameter choices was considered infeasible given the computational resources available for this study. Retrospectively, we see rather comparable performance levels across different input representations and involving independent code bases as an indication that the models/hyperparameter choices were not suboptimal.
Baseline Models. On VitalDB CalibFree, the baseline model (used for BP estimation) predicts BP by outputting the median SBP/DBP value of the BP data inferred from the training set for any given input, providing a straightforward reference for evaluating the performance of more advanced predictive models. On VitalDB Calib, we use The subject-specific median was calculated as the prediction on the test set.
Raw Time Series Models. For both BP estimation and AF detection, deep learning architectures such as CNNs, RNNs, and temporal convolutional networks (TCNs) have been used to process raw ECG or PPG sequences to capture complex temporal patterns, directly predicting continuous values for regression or output probabilities for classification. The specific models used in this study were:
@@
  • XResNet1d50/101: Deep residual networks modified for 1D data, incorporating group normalization and selective kernel sizes to learn hierarchical features from physiological signals [65].
  • XResNet1d50+GNNLL: This case uses the XResNet50d model with a Gaussian Negative Log Likelihood Loss (GNLL) instead of the conventional MAE loss as proposed in [38].
@@
  • PPNet: A hybrid architecture consisting of CNN and LSTM layers designed to extract spatial and temporal representations from physiological time series [53].
  • iTransformer: A transformer-based model tailored for time series prediction and analysis for instance-specific normalization techniques and attention mechanisms for improved sequence modelling [41].
  • TimesNet: A convolutional inception-like architecture for time series analysis which uses multi-periodic modelling by convolutional blocks for extracting detailed temporal patterns [75].
We can roughly categorize the considered models into complex/deep models (Inception1D, XResNet1d50, XResNet1d101, AlexNet1d, TCNs ) and simple/shallow models (LeNet1d, MiniRocket).
Feature-Based Models. Clinically interpretable features are extracted from PPG signals, such as pulse morphology metrics for BP estimation and irregularity measures for AF detection, and fed into machine learning algorithms to perform regression or classification tasks with clearer interpretability. The specific models and techniques used in this study were:
  • Clinically interpretable features (CIF): CIF for BP includes features such as systolic peaks, diastolic peaks, pulse duration, and pulse morphology metrics [9]. These features reflect vascular health, haemodynamic dynamics, and arterial compliance, making them well-suited for modelling BP. CIF for AF comprises features related to rhythm irregularity, such as randomness [13], variability [66, 39], and complexity in inter-beat-intervals [56]. These features highlight the irregular heart rhythms characteristic of AF. Table 3 provides an overview of the hand-crafted features used in the feature-based models.The full list of features is described in the supplementary material. These features are combined with the following prediction models:
@@
  • Wavelet Transformation: A Wavelet-based method such as Wavelet Packet Decomposition using the Discrete Wavelet Transform [58] is applied to the raw PPG signal. This transformation decomposes the signal into time-frequency components, extracting features that capture both transient and long-term patterns in the data. It should be noted that Wavelet + MLP forms a pipeline where the wavelet transform serves as a sophisticated feature extractor, and the MLP functions as the predictive model utilizing those features.
It should be noted that only the CIF directly reflects physiological mechanisms, e.g. pulse morphology or rhythm irregularity. On the other hand, features derived from mathematical transformations (e.g., wavelet coefficients) provide useful signal representations; however, they are not directly interpretable in clinical terms.
Table 3: Overview of feature categories and corresponding feature-based pipelines used for BP estimation and AF detection.
Pipeline Feature category BP AF Description / Examples
CIF Time-domain morphology Pulse amplitude, rise time, decay time, pulse duration, area under the pulse
CIF Derivative-based First and second derivative extrema, slope-based indices
CIF Temporal variability Inter-beat intervals (PP intervals), HR, HRV, rhythm irregularity measures
CIF Frequency-domain Spectral power in predefined frequency bands capturing rhythm variability
CIF Quality-related Signal quality indices, beat consistency measures
Wavelet + MLP Wavelet-based features Wavelet packet decomposition coefficients used as task-agnostic feature representations
Image-Based Models: One-dimensional signals can be converted into two-dimensional images, which capture the essence of the signal in a compact domain. Traditional image recognition CNNs can then be used with these image inputs. There are many different ways of converting signals to images, but this study only used the following two approaches:
  • Continuous Wavelet Transform (CWT) Scalograms: PPG signals are transformed into time-frequency images derived from PPG signals that capture localized signal variations [1]. The CWT-based images are used as inputs for ResNet18 models [28].
  • STFT-Based Spectrograms (Spectrogram-ResNet): PPG signals are transformed into fixed-resolution time-frequency spectrograms using the STFT [11]. The resulting spectrograms are processed using convolutional neural network architectures based on ResNet-50 and ResNet-18, initialized with ImageNet pre-trained weights [15, 29].
Table 10 lists overall parameter counts of all models within the tasks of AF detection and BP estimation, and is condensed by input representation (T/F/I). It is worth noting that parameter count is reported as a coarse, hardware-agnostic indicator of model complexity and is used in this work as a relative measure of computational efficiency, but should not be interpreted as a direct proxy for deployment cost on wearable or embedded hardware, where inference latency, memory footprint, and energy consumption are strongly platform dependent. For more details on these models and their implementations, please see the Supplementary Material.
@@ 4.1 BP Estimation
Tables LABEL:table_3a, LABEL:table_3b, LABEL:table_4a, and LABEL:table_4b display the results of BP (SBP/DBP) prediction using different deep learning models based on VitalDB Calib and VitalDB CalibFree datasets, respectively. The baseline models achieved MAEs of 14.87 and 9.43 mmHg for SBP and DBP, respectively, on CalibFree, and 10.72 and 5.78 mmHg, respectively, on Calib. On Calib, the per-subject baseline performed substantially better than the global baseline (10.72 and 5.78 mmHg vs. 14.91 and 9.52 mmHg respectively).
@@
Model performance varied between models and tasks. On CalibFree, almost all models provided an improvement over baseline for both SBP and DBP (the only exception being the Inception1d SBP model). However, the level of improvement was moderate at best, with the lowest MASEs of 0.83 for both SBP and DBP achieved by XResNet1d50, indicating 17% reductions in MAE in comparison to baseline. This resulted in at best 26% of SBP estimates and 40% of DBP estimates falling into the top grade (i.e. errors $\leq$ 5 mmHg). On Calib, there was greater variation in model performances, with XResNet1d50+GNLL achieving the lowest MASEs of 0.73 and 0.87 for SBP and DBP, respectively, corresponding to 48% of SBP estimates and 64% of DBP estimates falling into the top grade. In contrast, several models performed worse than the per-subject baseline, with the worst-performing model, CIF+MLP, achieving MASEs of 1.27 and 1.54 for SBP and DBP, respectively. Indeed, for DBP estimation, only the XResNet1d50+GNLL model achieved an improvement over the subject-specific baseline, whereas all others performed worse than this baseline. Absolute performance in terms of MAE was better on Calib than CalibFree, as shown by lower MAEs on Calib than CalibFree for all models except the TCN+MLP SBP and DBP models, and the CIF+MLP DBP model. MLP models always performed worse than other models of the same type.
@@
Table 4: The SBP performance on the VitalDB Calib dataset for three input representations, T for raw time series, F for feature-based, and I for image-based models, next to B for the baseline model. The best-performing model is marked in bold and underlined. Models that do not perform statistically significantly worse than the best model are shown in bold. All MAE values are given in units of mmHg. Bias and LoA were computed following Bland–Altman analysis (bias $\pm$ 1.96 SD).
Model SBP IEEE Grades (SBP)
MAE (MASE) Bias LoA A B C D
B Baseline (global) 14.91 (1.39) 0.00 36.94 0.21 0.04 0.04 0.71
Baseline (per subject) 10.72 (1.00) 0.00 28.01 0.34 0.06 0.05 0.55
T XResNet1d101 9.06 (0.84) 0.25 24.51 0.40 0.06 0.06 0.48
XResNet1d50 9.52 (0.88) 0.29 25.00 0.37 0.06 0.06 0.51
Inception1d 9.43 (0.88) -0.90 24.55 0.36 0.06 0.06 0.52
LeNet1d 11.76 (1.09) -0.78 29.40 0.27 0.05 0.05 0.63
XResNet1d50 +GNLL 7.99 (0.74) -0.79 22.44 0.48 0.06 0.05 0.41
AlexNet1d 9.49 (0.88) -0.54 25.21 0.38 0.06 0.06 0.50
Minirocket 11.24 (1.05) 0.11 28.39 0.29 0.06 0.05 0.60
iTransformer 12.66 (1.18) -0.35 31.71 0.25 0.05 0.05 0.65
TimesNet 12.69 (1.16) -0.71 31.88 0.25 0.05 0.05 0.65
PPNet 8.76 (0.81) -0.42 23.63 0.41 0.06 0.06 0.47
TCN +MLP 12.84 (1.19) -5.50 30.71 0.25 0.05 0.05 0.65
F WAVELET +MLP 13.62 (1.26) -1.77 34.10 0.24 0.05 0.04 0.67
CIF +GPR 12.22 (1.13) -0.32 32.03 0.27 0.05 0.05 0.63
CIF +MLP 13.78 (1.27) -0.13 34.21 0.23 0.04 0.04 0.69
I CWT 10.91 (1.01) -0.24 28.51 0.32 0.06 0.05 0.57
Spectrogram-ResNet-18 9.85 (0.91) -1.14 26.83 0.38 0.06 0.06 0.50
Spectrogram-ResNet-50 9.82 (0.91) -0.81 26.88 0.39 0.06 0.06 0.49
Table 5: The DBP performance on the VitalDB Calib dataset for three input representations, T for raw time series, F for feature-based, and I for image-based models, next to B for the baseline model. The best-performing model is marked in bold and underlined. Models that do not perform statistically significantly worse than the best model are shown in bold. All MAE values are given in units of mmHg. Bias and LoA were computed following Bland–Altman analysis (bias $\pm$ 1.96 SD).
Model DBP IEEE Grades (DBP)
MAE (MASE) Bias LoA A B C D
B Baseline (global) 9.52 (1.65) 0.00 23.66 0.32 0.06 0.06 0.56
Baseline (per subject) 5.78 (1.00) 0.00 15.37 0.56 0.07 0.06 0.30
T XResNet1d101 5.91 (1.02) -0.06 15.86 0.55 0.07 0.06 0.32
XResNet1d50 6.33 (1.08) 0.02 16.51 0.50 0.08 0.07 0.35
Inception1d 6.37 (1.11) -0.02 16.46 0.50 0.07 0.07 0.36
LeNet1d 7.80 (1.35) 1.00 19.34 0.39 0.07 0.07 0.47
XResNet1d50 +GNLL 5.09 (0.87) -0.52 14.56 0.64 0.06 0.05 0.25
AlexNet1d 6.10 (1.05) -0.14 16.33 0.54 0.07 0.06 0.33
Minirocket 7.33 (1.26) 0.01 18.51 0.43 0.07 0.07 0.43
iTransformer 8.20 (1.41) -0.18 20.53 0.38 0.07 0.06 0.49
TimesNet 8.12 (1.40) 0.01 20.69 0.38 0.07 0.06 0.49
PPNet 5.58 (1.01) -0.18 15.23 0.58 0.07 0.06 0.29
TCN +MLP 8.48 (1.46) -3.12 20.46 0.37 0.07 0.07 0.49
F WAVELET +MLP 8.84 (1.51) -1.00 22.08 0.36 0.07 0.06 0.51
CIF +GPR 7.78 (1.35) -0.10 19.69 0.40 0.07 0.07 0.46
CIF +MLP 8.94 (1.54) -0.55 22.18 0.35 0.06 0.06 0.53
I CWT 6.68 (1.15) -0.32 17.66 0.50 0.07 0.06 0.37
Spectrogram-ResNet-18 6.33 (1.09) -0.64 17.21 0.54 0.07 0.06 0.33
Spectrogram-ResNet-50 6.39 (1.10) -0.10 17.42 0.53 0.07 0.06 0.34
Table 6: The SBP performance on the VitalDB CalibFree dataset for three input representations, T for raw time series, F for feature-based, and I for image-based models, next to B for the baseline model. The best-performing model is marked in bold and underlined. Models that do not perform statistically significantly worse than the best model are shown in bold. All MAE values are given in units of mmHg. Bias and LoA were computed following Bland–Altman analysis (bias $\pm$ 1.96 SD).
Model SBP IEEE Grades (SBP)
MAE (MASE) Bias LoA A B C D
B Baseline 14.87 (1.00) -0.00 36.94 0.21 0.04 0.04 0.71
T XResNet1d101 12.43 (0.83) -1.28 30.91 0.25 0.05 0.05 0.65
XResNet1d50 12.46 (0.83) -0.23 31.09 0.24 0.05 0.05 0.66
Inception1d 14.46 (0.97) -3.15 30.62 0.25 0.05 0.05 0.65
LeNet1d 12.37 (0.83) 0.86 30.57 0.25 0.05 0.05 0.65
XResNet1d50+GNLL 12.27 (0.82) -0.99 30.55 0.26 0.05 0.05 0.64
AlexNet1d 12.51 (0.84) 0.52 31.24 0.25 0.05 0.05 0.65
Minirocket 12.65 (0.85) -0.95 31.43 0.26 0.05 0.05 0.64
iTransformer 13.15 (0.88) -0.75 32.76 0.24 0.05 0.05 0.66
TimesNet 13.09 (0.88) -1.89 32.84 0.25 0.05 0.05 0.65
PPNet 12.56 (0.84) -2.05 31.24 0.25 0.05 0.05 0.65
TCN +MLP 12.72 (0.85) -2.52 31.77 0.25 0.05 0.05 0.65
F WAVELET +MLP 14.21 (0.95) 0.03 35.32 0.22 0.04 0.04 0.70
CIF +GPR 12.90 (0.87) -0.32 32.38 0.25 0.05 0.04 0.66
CIF +MLP 14.02 (0.94) 0.15 34.82 0.24 0.04 0.04 0.68
CWT 13.49 (0.90) -1.14 33.63 0.23 0.05 0.04 0.68
Spectrogram-ResNet-18 13.01 (0.87) 0.32 32.29 0.24 0.04 0.04 0.68
I Spectrogram-ResNet-50 13.53 (0.90) 3.40 32.88 0.24 0.04 0.04 0.68
Table 7: The DBP performance on the VitalDB CalibFree dataset for three input representations, T for raw time series, F for feature-based, and I for image-based models, next to B for the baseline model. The best-performing model is marked in bold and underlined. Models that do not perform statistically significantly worse than the best model are shown in bold. All MAE values are given in units of mmHg. Bias and LoA were computed following Bland–Altman analysis (bias $\pm$ 1.96 SD).
Model DBP IEEE Grades (DBP)
MAE (MASE) Bias LoA A B C D
B Baseline 9.43 (1.00) 0.00 23.66 0.33 0.06 0.06 0.55
T XResNet1d101 7.93 (0.84) 0.18 19.70 0.39 0.07 0.06 0.48
XResNet1d50 7.98 (0.84) -1.55 19.63 0.39 0.07 0.06 0.48
Inception1d 7.94 (0.84) -1.86 19.29 0.38 0.07 0.06 0.49
LeNet1d 7.92 (0.84) 1.26 19.45 0.39 0.07 0.06 0.48
XResNet1d50+GNLL 7.94 (0.84) -0.54 19.63 0.39 0.07 0.06 0.48
AlexNet1d 7.98 (0.84) 0.15 19.80 0.38 0.07 0.07 0.48
Minirocket 8.08 (0.85) -1.01 19.87 0.38 0.07 0.06 0.49
iTransformer 8.37 (0.88) -0.67 20.77 0.37 0.07 0.06 0.50
TimesNet 8.29 (0.87) -0.71 20.79 0.38 0.07 0.06 0.49
PPNet 7.95 (0.84) -0.91 19.74 0.39 0.07 0.07 0.48
TCN +MLP 8.24 (0.87) -1.77 20.27 0.38 0.06 0.06 0.50
F WAVELET +MLP 8.89 (0.94) -0.78 22.10 0.35 0.06 0.06 0.53
CIF +GPR 8.15 (0.86) -0.44 20.41 0.38 0.07 0.06 0.49
CIF +MLP 8.69 (0.92) 0.38 21.61 0.36 0.06 0.06 0.52
CWT 8.41 (0.89) -0.18 20.96 0.37 0.07 0.07 0.49
I Spectrogram-ResNet-18 8.04 (0.85) 0.27 19.92 0.38 0.07 0.07 0.48
Spectrogram-ResNet-50 8.34 (0.88) 2.00 20.30 0.37 0.07 0.06 0.50
@@
Raw time series (T) and image-based (I) models performed better than feature-based (F) models on Calib (with the one exception of TCN+MLP), as shown by MASEs of 0.73-1.07 (SBP) and 0.87-1.46 (DBP) for raw time series and image-based models, compared to 1.13-1.27 and 1.35-1.54 for feature-based models. There was a less clear difference in performance on CalibFree. Whilst best performance was achieved with raw time series models, the image-based model performed better than three raw time series models on Calib, and better than one raw time series model on CalibFree. For Calib, more complex models seem to exhibit an advantage (comparing, for example, XResNet1d50 and XResNet1d101), most likely due to the ability to memorize subject-specific signal patterns. The only exception from this pattern seems to be the AlexNet1d model, which shows a performance that is almost on par with more complex models and considerably better than comparable lightweight models (such as LeNet1d or TCN+MLP).
@@
Absolute performance in terms of MAE was always better for DBP estimation than for SBP estimation. However, when considering errors relative to baseline, MASEs were broadly similar between SBP and DBP on CalibFree, and always higher for DBP than SBP on Calib. Since the IEEE grading system uses absolute errors, performance in terms of IEEE Grades was generally better for DBP than SBP.
4.2 AF Detection
Table 8 represents the performance analysis of the classification task (AF/ non-AF) based on the Deepbeat dataset. Specificity (sensitivity $\geq 0.8$) and sensitivity (specificity $\geq 0.8$) are reported at operating thresholds selected to ensure a minimum sensitivity or specificity of 0.8, respectively.
Similar to the regression task, raw time series (T) and image-based (I) models performed better than feature-based (F) models in all cases except for the LeNet1d raw time series model, which performed poorly. The best-performing models were Inception1d and XResNet1d50, which achieved AUCs of 0.87, and F1 scores of 0.72 and 0.69 respectively. As with the regression task, whilst best performance was achieved with the more complex raw time series models, the image-based model performed better than some raw time series models. Whilst feature-based models generally performed worst, wavelet-based features produced better performance than clinically interpretable features.